What is the real solutions of the equation 3(x+1)^2+16(x+1)+5=0
Answer by robertb(5830)

(3(x+1) + 1)((x+1) + 5) = 0 by direct factoring of the left side.
<==> (3x + 3 + 1)(x + 1 + 5) = 0
<==> (3x + 4)(x + 6) = 0
==> x = -4/3, -6.